# Visatronic: A Multimodal Decoder-Only Model for Speech Synthesis

### v1

@@ 4 Related Work
Text-to-Speech Synthesis Text-to-speech (TTS) systems have evolved from early approaches to end-to-end methods [49, 16, 24, 27, 31, 34, 35]. Traditional TTS systems face significant challenges with unseen speaker styles due to substantial enrollment data requirements. While several approaches attempt to address this by extracting speaker representations from speech data [6, 14, 15, 22, 28], obtaining sufficient high-quality utterances remains problematic. Recent studies have incorporated face images for speaker representation [12, 21, 42], aiming to capture visual-acoustic correlations, but often neglect motion-related factors leading to inconsistent voice generation when facial expressions vary. Recent unified architectures for speech-text modeling like VioLA [43] require multi-stage hierarchical processing for EnCodec [11] features, while VOXTLM [25] uses an LM-style approach but relies on HuBERT content tokens losing acoustic and speaker characteristics.
@@
Lip-to-Speech Synthesis Lip-to-speech synthesis aims to reconstruct speech signals from a given face image and silent videos of lips of talking-face, crucial for scenarios with corrupted or missing audio. Early approaches used encoder-decoder architectures with GAN-based training – Lip2Wav [32], End-to-end GAN [30], and VCA-GAN [17] demonstrated success on limited vocabulary datasets, while Lip2Speech [18] extended the GAN framework with multi-task learning for improved content modeling. Recent advances explored discrete token representations through AV-HuBERT [36], with works like ReVISE [13] integrating HiFi-GAN for improved audio generation. In parallel, diffusion models have emerged as a powerful approach for speech generation. Works like DiffWave [20], Grad-TTS [31], and PriorGrad [23] demonstrated effective speech synthesis, leading to LipVoicer [46] which adapted diffusion models for lip-to-speech generation. However, these approaches focus primarily on lip movements, potentially overlooking broader visual dynamics that could improve speech generation. Our work takes a different direction by proposing a novel video-text-to-speech task that leverages complete visual context alongside text input. Rather than using GAN or diffusion-based approaches, we adopt a unified decoder-only transformer architecture inspired by recent successes in LLMs. This enables seamless integration of video, text, and speech modalities for more natural and contextually appropriate speech generation.
@@ 5 Conclusion
To the best of our knowledge, we are the first to propose a video-text-to-speech generation framework using a decoder-only transformer architecture. This approach simplifies the multimodal conditional generation of speech while maintaining high-quality output. We demonstrate our approach’s effectiveness by achieving state-of-the-art performance on two challenging datasets: VoxCeleb2 and LRS3 compared to prior approaches that used cropped-lip inputs. These datasets feature diverse speakers, accents, and recording conditions, showcasing our model’s ability to handle real-world scenarios. We formulated a suite of evaluation metrics including Mean Opinion Score for style, synchronization, and content to evaluate naturalness and overall quality of the generated speech. In addition we proposed an automatic metric to assess the quality of alignment between generated and original speech. This multi-faceted evaluation goes beyond traditional metrics to capture nuanced aspects of speech synthesis quality.
@@ Appendix D Video Reconstruction
Appendix E Evaluation Metrics
@@
SyncScore   We are using open-source code for SyncNet from https://github.com/joonson/syncnet_python. During evaluation if generated speech is longer than the video duration then the last video frame is used for the rest of the speech duration.
TimeSync   We use https://github.com/richardbaihe/a3t from [2] to do force alignment between the phoneme sequence of ground truth transcription (PL.v2 is treated as a ground truth for VoxCeleb2) and speech: either generated or original audio. The code is using HMM model from HTK [47] to perform force alignment. This procedure gives us phoneme location in time and phoneme duration for each audio. After, we exclude the silence (“sp”) and its duration from each alignment, see Figure 8. Because every word has several possible phoneme sequences, we use Levenstein distance computation to align sequence of phonemes obtained for the generated and original audio: we consider phonemes to be aligned if they are equal or they can be obtained via substitution operation. Then, we compute the average absolute time difference between the centers of each aligned phoneme segments in generated and original audio, see Figure 9.
TimeSync can be expressed as $\frac{1}{N}\sum_{\phi^{GT}}|t^{model}_{f(\phi^{GT})}-t^{GT}_{\phi^{GT}}|$, where $N$ is the total number of phonemes $\phi^{GT}$ in the ground truth transcriptions obtained for the original audio samples; $t^{GT}_{\phi^{GT}}$ is the segment’s center (in seconds) for the phoneme $\phi^{GT}$ and $t^{GT}_{\phi^{GT}}=(start_{\phi^{GT}}+end_{\phi^{GT}})/2$; $f(\phi^{GT})$ is the phoneme in the phoneme sequence of the generated audio sample which corresponds to ${\phi^{GT}}$ in the alignment; $t^{model}_{f(\phi^{GT})}$ is the segment’s center (in seconds) for the phoneme $f(\phi^{GT})$ and $t^{model}_{f(\phi^{GT})}=(start_{f(\phi^{GT})}+end_{f(\phi^{GT})})/2$; $start$ and $end$ indicate the phoneme’s start and end timestamps in generated or ground truth audio.
Figure 8: TimeSync. Example of the phoneme sequence and its alignment for the ground truth audio before removing silence (“sp”) segments (blue) and after (green).
Figure 9: TimeSync. Example of the phoneme sequence for the ground truth audio (green) and corresponding aligned phoneme sequence for the generated audio (red). TimeSync is computed on these paired (green and red) segments by taking the absolute difference between the segments centers (measure in seconds).
@@
We further instruct raters to evaluate emotional consistency between video and generated speech (’video-speech emotions’) and emotional expressiveness in speech (’speech emotions’) by comparing ground truth and generated audios, see instructions in Figures 11 and 12. MOS results, Table 8, demonstrate the benefit of visual conditioning for emotional expressiveness.
Method video-speech emotions ($\uparrow$) speech emotions ($\uparrow$)
GT 4.62 ±0.07 4.92 ±0.04
@@
VTTS (VT-ordered) 3.74 ±0.12 3.39 ±0.15
Table 8: Human evaluation of speech-video emotion alignment on VoxCeleb2. Mean Opinion Scores (1-5) with 95% confidence intervals measuring emotional consistency between video and generated speech (’video-speech emotions’) and emotional expressiveness in speech (’speech emotions’). VTTS (TV-ordered) achieves the best performance in video-speech emotion alignment (3.79), while VTTS (VT-ordered) performs best in speech emotion quality (3.39). Both significantly outperform TTS baseline, demonstrating the benefit of visual conditioning for emotional expressiveness. Ground truth (GT) serves as reference, while GT (discrete) shows the upper bound achievable with our approach.
Appendix F Implementation Details
VoxCeleb2 original data has video at 25fps (40ms per frame), or 25Hz, which we use for video representation extraction, while the audio is given at 16kHz and we extract speech representations at 40Hz (25ms per frame).
@@ Appendix F Implementation Details
Appendix G Video-to-Speech
@@
Appendix H Qualitative Results
In Figures 1416, and 18, we show log mel-spectrogram comparisons between TTS, Ground Truth (GT), and our VTTS (VT-ordered) model across different scenarios. These visualizations include both successful cases where VTTS (VT-ordered) effectively captures temporal dynamics and spectral patterns, and failure cases (Figure  18) that highlight current limitations. Through these examples, we can analyze how video conditioning helps maintain proper speech duration and temporal alignment, while also identifying challenges in generating complex spectral information. Furthermore, to analyze temporal synchronization between generated and ground truth speech, we visualize phoneme-level alignments in Figures 14, 16, and 18. Each plot shows the relationship between phoneme timings in ground truth ($x$-axis) versus generated speech ($y$-axis), where perfect synchronization would follow the diagonal dashed line. The different variants of our model, VTTS (VT-ordered) consistently demonstrate better temporal alignment compared to TTS, as evidenced by their closer distance to the ideal diagonal. This visualization helps quantify how video conditioning helps to maintain proper speech timing and rhythm, with VTTS (VT-ordered) variants showing improved temporal coherence across different examples.

### v2

@@ 4 Related Work
Speech synthesis models have made rapid progress in recent years, yet they typically focus on text-to-speech or simplified variants of video-conditioned speech generation, e.g. from cropped lip videos [55], or generic audio generation [25]. While the former leverages pretrained lip detectors to simplify the problem, it omits important visual cues from the full face that contribute to accurate modeling of phoneme sounds particularly those involving regions beyond the lips. Furthermore, prior approaches often rely on pretrained ASR [55] or follow a pipeline of lip-reading followed by speech generation, which can degrade in noisy, multi-speaker environments where parts of the speech signal may be lost or masked by background interference. To this end, we tackle the video-text to speech, multimodal speech synthesis task, where the goal is to generate temporally aligned and intelligible speech from a speaker’s full-face video and transcript.
Text-to-Speech (TTS) Synthesis
@@
TTS systems have evolved from early approaches to end-to-end methods [58, 22, 30, 34, 38, 41, 43]. Traditional TTS systems face significant challenges with unseen speaker styles due to substantial enrollment data requirements. While several approaches attempt to address this by extracting speaker representations from speech data [7, 19, 20, 28, 35], obtaining sufficient high-quality utterances remains problematic. Recent studies have incorporated face images for speaker representation [16, 27, 50], aiming to capture visual-acoustic correlations, but often neglect motion-related factors leading to inconsistent voice generation when facial expressions vary. Recent unified architectures for speech-text modeling like VioLA [51] require multi-stage hierarchical processing for EnCodec [14] features, while VOXTLM [32] uses an LM-style approach but relies on HuBERT content tokens losing acoustic and speaker characteristics.
Lip-to-Speech (LTS) Synthesis
@@
LTS aims to reconstruct speech signals from a given face image and silent videos of lips of talking-face, crucial for scenarios with corrupted or missing audio. Early approaches used encoder-decoder architectures with GAN-based training – Lip2Wav [39], End-to-end GAN [37], and VCA-GAN [23] demonstrated success on limited vocabulary datasets, while Lip2Speech [24] extended the GAN framework with multi-task learning for improved content modeling. Recent advances explored discrete token representations through AV-HuBERT [44], with works like ReVISE [17] integrating HiFi-GAN for improved audio generation. In parallel, diffusion models have emerged as a powerful approach for speech generation. Works like DiffWave [26], Grad-TTS [38], and PriorGrad [29] demonstrated effective speech synthesis, leading to LipVoicer [55] which adapted diffusion models for LTS generation. However, these approaches focus primarily on lip movements, potentially overlooking broader visual dynamics that could improve speech generation. Our work takes a different direction by considering complete visual context alongside text input. Rather than using GAN or diffusion-based approaches, we adopt a unified decoder-only transformer architecture inspired by recent successes in LLMs. This enables seamless integration of video, text, and speech modalities for more natural and contextually appropriate speech generation.
Visual Voice Cloning (V2C) and Dubbing
Adjacent research, V2C [8, 21] and dubbing [11, 18, 59, 12, 31], address related challenges but operates under different setting than VTTS. V2C typically focuses on character-specific voice synthesis, aiming to clone a target speaker’s vocal identity and speaking style, often using a short reference audio-video pair without explicit transcript supervision or temporal alignment to facial motion. In contrast, dubbing replaces a video’s original speech with new content in a different voice or language, often relying on reference speech, prosody labels, or style embeddings to guide generation. While recent V2C extensions consider text-conditioned generation, most rely on lip crops and operate on constrained speaker sets (e.g., GRID [15], CHEM [2]), whereas our model uses full-face video, targets perceptual consistency instead of identity cloning, and generalizes to unseen speakers in unconstrained settings like VoxCeleb2 [10] and LRS3 [1]. Our model generates emotionally expressive and temporally aligned speech from video-text pairs, using speaker embeddings but no reference audio or handcrafted prosody supervision – thus framing expressive dubbing as an end-to-end generative modeling problem.
@@ 5 Conclusion
To the best of our knowledge, this is the first work to formalize and benchmark the video-text to speech (VTTS) task using a unified decoder-only transformer. Our approach simplifies multimodal speech generation by jointly modeling full-face video and transcripts without relying on lip crops, handcrafted prosody labels, or reference audio. We demonstrate strong generalization on two challenging benchmarks – VoxCeleb2 and LRS3 – which feature diverse speakers, accents, and unconstrained acoustic conditions. To evaluate the quality of generated speech, we employ both subjective and objective metrics, including mean opinion scores for intelligibility, naturalness, synchronization, and emotional expressiveness. Additionally, we introduce TimeSync, a phoneme-level alignment metric to assess temporal synchronization between generated and ground-truth speech. This comprehensive evaluation captures both content fidelity and expressive alignment, highlighting the effectiveness of our proposed framework.
@@ Appendix D Video Reconstruction
Appendix E Video Representation
To obtain a frame-level embedding suitable for autoregressive modeling, we aggregate the spatial grid of quantized video embeddings into a single vector $\mathbf{z}_{t}^{v}\in\mathbb{R}^{D^{\prime}}$ per frame. As shown in Figure 2 (main paper), each frame is first discretized using a pretrained VQ-VAE encoder, resulting in a $H^{\prime}\times W^{\prime}$ grid of tokens, where each token is embedded via a learnable table $\mathbf{E}^{v}$. Since our unified decoder processes one token per time step, we aggregate the $H^{\prime}\times W^{\prime}$ spatial embeddings into a single vector before feeding into the decoder model. Below, we outline the aggregation strategies we explored:
Max pooling: $\mathbf{z}_{t}^{v}=\max_{(h,w)}\mathbf{e}^{v}_{t,h,w}$;
Stacking: stack embeddings and then project it via a learnable linear layer $\mathbf{L}^{v}(\cdot):\mathbb{R}^{H^{\prime}W^{\prime}D^{\prime}}\to\mathbb{R}^{D^{\prime}}$, $\mathbf{z}_{t}^{v}=\mathbf{L}^{v}([\mathbf{e}^{v}_{t,1,1},\mathbf{e}^{v}_{t,1,2},\dots,\mathbf{e}^{v}_{t,1,W^{\prime}},\mathbf{e}^{v}_{t,2,1},\dots,\mathbf{e}^{v}_{t,H^{\prime},W^{\prime}}])$.
Appendix F Evaluation Metrics
@@
We further instruct raters to evaluate emotional consistency between video and generated speech (’video-speech emotions’) and emotional expressiveness in speech (’speech emotions’) by comparing ground truth and generated audios, see instructions in Figures 8 and 9. MOS results, Table 67 highlight the advantages of visual conditioning for intelligibility, naturalness, synchronization, and emotional expressiveness. We use emotions2vec [^3] https://huggingface.co/emotion2vec to select audio with non-neutral emotions (6 classes, each prob. $>99\%$).
Table 6: Human evaluation on VoxCeleb2. Mean Opinion Scores (MOS) (1-5) with 95% confidence intervals for Intelligibility, Naturalness, and Synchronization. VTTS (VT-ordered) achieves the best performance in Intelligibility (3.48) and Naturalness (3.20), while VTTS (TV-ordered) performs best in Synchronization (2.50). Both significantly outperform TTS baseline across all metrics. Ground truth (GT) serves as a reference, and GT (discrete) is an upper bound due to speech discretization.
GT 4.55 ±0.09 4.79 ±0.05 4.57 ±0.10
GT (discrete) 3.95 ±0.13 3.77 ±0.15 4.36 ±0.12
TTS 3.17 ±0.19 2.92 ±0.21 1.98 ±0.15
VTTS (TV-streaming) 3.19 ±0.17 2.99 ±0.16 2.28 ±0.17
VTTS (TV-ordered) 3.35 ±0.17 3.02 ±0.19 2.50 ±0.21
VTTS (VT-ordered) 3.48 ±0.15 3.20 ±0.19 2.48 ±0.19
Table 7: Human evaluation of emotional expressiveness on VoxCeleb2. Mean Opinion Scores (1-5) with 95% confidence intervals for emotional consistency between video and speech (video-speech emotions), perceived emotion quality in speech alone (speech emotions), and specific emotion categories (Angry, Happy, Fearful). VTTS (TV-ordered) achieves the highest score in video-speech emotion alignment (3.79), while VTTS (VT-ordered) performs best in overall speech emotion quality (3.39) and in all individual emotion categories. These results demonstrate the advantage of visual conditioning for expressive speech generation. Ground truth (GT) serves as the reference, and GT (discrete) provides the upper bound achievable with our quantization-based approach.
Method video-speech emotions ($\uparrow$) speech emotions ($\uparrow$) Angry ($\uparrow$) Happy ($\uparrow$) Fearful ($\uparrow$)
GT 4.62 ±0.07 4.92 ±0.04 4.36 4.43 4.20
@@
VTTS (VT-ordered) 3.74 ±0.12 3.39 ±0.15 3.82 3.76 3.73
Appendix G Implementation Details
Datasets    1) LRS3 [1] is audio-visual dataset in English, compiled from TED and TEDx video presentations. This dataset stands out for its focus on unconstrained long sentences, featuring a rich vocabulary of over 50k words and thousands of unique speakers. It contains approximately 151k videos with around 439h of speech with transcription. There are 1,452 videos in the test split. 2) VoxCeleb2 [10] is a large-scale audio-visual dataset primarily designed for speaker recognition task but applicable to various audio-visual processing domains. It consists of over 1M face-cropped YouTube videos from more than 6k distinct identities, resulting in 1.6k hours of speech w/o paired transcription. The dataset is characterized by high variability in lighting conditions, image quality, pose, and motion blur, with an average video duration of 8s. This diversity in real-world conditions makes VoxCeleb2 particularly useful for developing robust models capable of performing well in unconstrained environments. Original data has video at 25fps (40ms per frame), or 25Hz, which we use for video representation extraction, while the audio is given at 16kHz and we extract speech representations at 40Hz (25ms per frame).
Hyper-Parameters Tuning
@@ Appendix G Implementation Details
Appendix H Ablation
Faster Convergence
Table 8 shows results for different number of training steps. Our models achieve better performance at 2M iterations compared to TTS baseline, showing that video modality is speeding up training convergence.
Table 8: Convergence analysis.   Training iterations comparison shows faster convergence when video modality is used in addition to text. While TTS requires 3M iterations, both VTTS (VT-ordered) and VTTS (TV-ordered) variants achieve comparable or better performance in only 2M iterations. LTTS (LT-ordered) is also need more training steps to converge compared to VTTS (VT-ordered). VTTS (VT-ordered) demonstrates superior performance with 12.2% WER and smaller gap (+2.1%) from GT (discrete) WER, showing efficient optimization when leveraging both modalities.
TTS 2M 2.6 ±0.1 10.1 ±0.2 17.3$\mathord{+}7.2$
VTTS (TV-ordered) 17.0$\mathord{+}6.9$
LTTS (LT-ordered) 14.9$\mathord{+}3.9$
VTTS (VT-ordered) $\mathord{+}2.1$
TTS 3M 2.6 ±0.1 10.1 ±0.2 14.7$\mathord{+}4.6$
VTTS (TV-ordered) 14.1$\mathord{+}4.0$
LTTS (LT-ordered) 14.0$\mathord{+}3.9$
VTTS (VT-ordered) $\mathord{+}2.1$
TimeSync metric for LTTS voxceleb
We run the TimeSync metric for LTTS and got 0.46 ±0.61, whereas the full video (VTTS, TV-ordered) gives 0.44 ±0.65, showing full-face video helps instead of lip-only crops. This indicates that the model benefits from additional visual cues beyond the lips which contribute to better temporal alignment between the generated and ground-truth speech.
LRS3 WER Computation
Interestingly, our VTTS models trained solely on VoxCeleb2 outperform those trained directly on LRS3, despite being evaluated on LRS3. We attribute this to the limited scale and diversity of LRS3, which contains fewer speakers and more constrained speaking conditions. In contrast, VoxCeleb2 offers a broader distribution of speaker identities, accents, and environmental variability, enabling more robust cross-modal alignment. This highlights the value of large, diverse pretraining even for in-domain tasks with smaller datasets.
Table 9: Objective evaluation on LRS3. Word error rate (WER) results for various VTTS models and baselines. Models trained on VoxCeleb2 generalize surprisingly well to LRS3 in zero-shot settings, outperforming prior methods trained directly on LRS3 (denoted by †).
Method Training Dataset WER (%) $\downarrow$
LipVoicer† [55] LRS3 21.4
VTTS (TV-ordered, zero-shot) VoxCeleb2 4.5
VTTS (VT-ordered, zero-shot) VoxCeleb2 8.2
VTTS (TV-ordered) LRS3 6.3
Appendix I Video-to-Speech
@@
Appendix J Qualitative Results
In Figures 1113, and 15, we show log mel-spectrogram comparisons between TTS, Ground Truth (GT), and our VTTS (VT-ordered) model across different scenarios. These visualizations include both successful cases where VTTS (VT-ordered) effectively captures temporal dynamics and spectral patterns, and failure cases (Figure  15) that highlight current limitations. Through these examples, we can analyze how video conditioning helps maintain proper speech duration and temporal alignment, while also identifying challenges in generating complex spectral information. Furthermore, to analyze temporal synchronization between generated and ground truth speech, we visualize phoneme-level alignments in Figures 11, 13, and 15. Each plot shows the relationship between phoneme timings in ground truth ($x$-axis) versus generated speech ($y$-axis), where perfect synchronization would follow the diagonal dashed line. The different variants of our model, VTTS (VT-ordered) consistently demonstrate better temporal alignment compared to TTS, as evidenced by their closer distance to the ideal diagonal. This visualization helps quantify how video conditioning helps to maintain proper speech timing and rhythm, with VTTS (VT-ordered) variants showing improved temporal coherence across different examples. Analysis of TimeSync for synchronization is given in Figure 16 .